A golfer played 8 rounds on a tournament course with the following scores (par is the expected score for a round; negative numbers represent the number of strokes under par for the round and positive numbers represent the numbers of strokes over par for the round): +5, -2, +3, -1, -3, +6, +2, -7. The golfer played 2 more rounds, ending with an even par total for the 10 rounds. Her score for the last was 7 more than her score on the next-to-last round. What scores did the golfer have on the last two rounds?
Answer by rwm(914)

so far on the 8 rounds the golfer is 3 over.
x+7+x+3=0
2x+10=0
2x=-10
x=-5
x+7=2
-5 and 2 are the final two rounds